OpenStudy (anonymous):

How can you find the exact value of sec(315), as well as csc(60)?

OpenStudy (amoodarya):

cos(315)=cos(360-45) =cos(-45)=cos(45) sec(315)=1/cos(315)=1/cos(-45)=1/cos(45)=square root(2)